Find a general form of an equation for the perpendicular bisector of the segment AB. A(4,-4),B(-2,4)
Answer by mananth(16946) 
 x1 y1 x2 y2
4 -4 -2 4
slope m = (y2-y1)/(x2-x1)
( 4 - -4 )/( -2 - 4 )
( 8 / -6 )
m= -4/3
The slope of a perpendicular line will be the negative reciprocal of this slope
=3/4
slope = 3/4 and it passes through the midpoint of the above two points
mid point formula
x= (x1+x2)/2 = (4-2)/2=1
y=0
the point is (1,0)
slope = 3/4, point = (1,0)
substitute in y=mx+b
0=3/4 *1+b
b=-3/4
the equation is y=(3/4)x-3/4
|
|
|
| |